OpenStudy (anonymous):

a 230.0ml sample of a 0.275 molarity solution is left on a hot plate over night the folowing morning the solution is 1.5 molarity. what volume of solvent has evaportated from the 0.275 molarity solution?

OpenStudy (anonymous):

Mi=n/Vi, Mf=n/Vf => n=Mi*Vi => Vf=Mi*Vi/Mf where M are molarities(initial and final with i and f), V the volumes, n the moles So you have Vf=0.275*230/1.5=42.17 ml ;)